# Micellization in active matter of asymmetric self-propelled particles: Experiments

### v2

@@ Abstract
Active matter composed of self-propelled particles features fascinating self-organization phenomena, spanning from motility-induced phase separation to phototaxis to topological excitations depending on the nature and parameters of the system. In the present paper, we consider micelle formation by active particles with a broken symmetry having a circular back and a sharpened nose toward which the particles accelerate. As we demonstrate in experiments with robotic swarms, such particles can either remain in the isotropic phase or form micelles depending on the location of their center of inertia, in accordance with a recent theoretical proposal [T. Kruglov, A. Borisov, Particles 2021, 2 (2021)]. Such a behavior is observed for both non-chiral particles moving linearly and placed in a parabolic potential and for chiral particles moving along circular trajectories on a flat surface. By performing experiments with single robots and two-robot collisions, we unveil that the observed emergence of micellization associated with shifting robots’ center of inertia towards their noses is governed by at least two-particle effects, in particular, by a difference in the formation of stable two-robot clusters. Finally, we consider the dependence of micelle lifetime and formation probability as well as two-robot collisions on friction between the lateral surfaces of the robots. Crucially, the predicted micellization does not involve any solvation shells that give rise to the micellization of surfactants but is instead driven by an interplay of activity and particle shape asymmetry.
@@ I Introduction
However, the formation of micelles – round or spherical-shaped assemblies of elongated particles in which they orient one of their non-equivalent edges to the inner region of a cluster and the other one to the outer region – has been demonstrated only in Janus particles covered with surfactants Zhang et al. (2015). This phenomenon seemed elusive in self-propelled particles without charge displacement, until the recent theoretical proposal of active matter micellization driven purely by particle shape asymmetry Kruglov and Borisov (2021) and experimental studies of rotelles, the rotational analogs of micelles formed by surfactant-like chain of self-rotating active particles Scholz et al. (2021).
@@
In the present paper, we consider ensembles of self-propelled particles that have the teardrop shape shown in Fig. 1(a). Particles having a similar shape occur in biological systems, for example, specific types of blood cells Gütgemann et al. (2014), and can be fabricated on the millimeter to micrometer scale with the help of various techniques, e.g. self-propelled asymmetric Janus particles in Refs. Dai et al. (2016); Shah et al. (2021); Liu et al. (2022). To demonstrate experimentally that such particles can form micelle-like clusters depending on the center of inertia location at the particle axis, we construct a swarm of teardrop-shaped vibrating robots (bristle-bots) shown in Fig. 1(b),(c) which are based on the Swarmodroid 1.0 platform Dmitriev et al. (2023). Such robots move with a controlled level of activity and are placed in a shallow parabolic potential of a satellite dish to prevent condensation at the system boundary, typical of self-propelled particles Giomi et al. (2013); Deblais et al. (2018); Boudet et al. (2021), Fig. 1(a), or on a flat surface with their trajectories modified to be slightly chiral (corresponding to robots moving along circular loops). In accordance with the predictions of Ref. Kruglov and Borisov (2021), we report micelle formation both for swarms on a flat surface and in a parabolic potential and address the dependence of their formation probability on the packing density of robots, their center of inertia location, velocity of their motion, and the friction coefficient between lateral surfaces of robots. Finally, we uncover the microscopic mechanisms responsible for such a sharp dependence of the micellization on the location of individual robots’ center of inertia by performing experiments with single robots as well as with two-robot collisions in different conditions.
@@
The paper is organized as follows. In Section II, we describe our experimental platform. Then, in Section III a key quantitative parameter is introduced which we use to detect micellization. Section IV presents the results of experiments with robotic swarms and includes three subsections. Subsection IV.1 includes studies of the condensation of robots at the boundary when they propel along straight trajectories on a flat surface. The next Subsection IV.2 features a demonstration of robots’ micellization on a flat surface for robots moving along chiral trajectories and includes the study of the micellization statistical properties. Then, Subsection IV.3 considers micellization of robots in a parabolic potential, including the dependence of the lifetime and formation probability of micelles on the packing density of robots, their motion velocity, and the friction between their lateral surfaces for different locations of the robots’ center of inertia. The experiments considering microscopic origin of the observed micellization transition are discussed in Section V. The first Subsection V.1 includes studies of individual robot motion in a parabolic potential for different center of inertia locations demonstrating that the behavior at a single-robot level does not demonstrate any pronounced changes. The following Subsection V.2 features the studies of two-robot collisions pointing toward a considerable increase in the formation of stable quarter-micelle clusters, hinting towards its possible influence on the emergence of the micellization. Section VI contains the final remarks and discussion of the obtained results.
@@ II Experimental platform for active matter: asymmetric robots
In our experiments, we implement the teardrop-shaped particles as self-propelled bristle-bots converting their vibration to a directed motion, Fig. 1(b)-(e). Such robots consist of a printed circuit board [Fig. 1(d)] carrying a vibration motor, a battery, and circuitry for infrared remote control, which allows turning the robots on and off simultaneously and varying their vibration activity (i.e., a self-propulsion velocity) by means of pulse width modulation (PWM) of the motor voltage, and a 3D-printed body with elastic bristles at the bottom [Fig. 1(e)] Dmitriev et al. (2023). The system dynamics is extracted using a tracking pipeline based on our custom recognition software111https://github.com/swarmtronics/ampy and the use of ArUco markers.
The length of each robot is , the diameter of the round part is , and the nose angle is corresponding to particles in a complete micelle. We choose such a value of to clearly distinguish micellization from crystallization, as the point symmetry is incompatible with a crystalline order. The height of the robots including the bristles is mm. The robot’s center of inertia lies on its mirror symmetry axis at the distance from its nose cusp. With the help of an additional load consisting of an M616 screw and three M6 nuts made of steel, which are fastened to a hole near the robot’s nose, we are able to move the center of inertia from mm (point ) corresponding to the absence of additional load [Fig. 1(b)] to mm (point ) for a robot with a load [Fig. 1(c)]. In the following, we denote robots with and without the additional load as -robots and -robots, respectively. The -robot mass is g and the -robot mass is g. We study swarms composed of , , , and robots.
@@ III Micellization parameter
where and are the in-plane coordinates of the noses of particles with indices and , is the number of particles in a complete micelle (in our case, ), is the characteristic distance between the noses of particles at which they are considered forming a micelle (here, we choose mm), and is the total number of particles in the system. Figure 2(a) shows the values of such micellization parameter extracted experimentally for several model systems: an isolated complete micelle with , several types of isolated incomplete micelles with , an inverse micelle with , a structure of densely packed robots with crystalline order, a single robot, and two types of two-robot clusters (“tandem” and “heart”) that, as will be shown in Section V, are important for understanding the processes facilitating micelle formation. The terms in the sum by index tend to unity for a complete micelle while rendering values orders of magnitude smaller for other configurations. However, due to the exponential dependence on the distance between the robots’ noses, the micellization parameter yields values somewhat lower than unity for experimentally observed micelles, with the majority of observed micelles corresponding to . For incomplete micelles and non-micellar structures, the micellization parameter yields exponentially low values. The normalization prefactor ensures the invariance of the parameter with respect to the number of robots. For fully chaotic systems, the micellization parameter yields the value independently of the number of robots in the system. The same value corresponds to the system consisting of a single robot and is the lowest possible value of the micellization parameter.
@@
We also study the dependence of the micellization parameter on the distance between the particles in a single micelle by introducing random shifts of the particles’ noses from the center of a micelle, Fig. 2(b). The components of the shift vectors are normally distributed with the zero mean and the dispersion that varies from mm to mm (approximately of the robot’s diameter). The micellization parameter decreases exponentially depending on this amplitude, producing the same value for a complete micelle with mm gaps between the robots and an ideal incomplete micelle without one robot.
IV Experimental studies of micellization in swarms of robots
We start with considering a swarm of teardrop-shaped robots moving on a flat surface and confined by a circular barrier, Table 1. The barrier is 3D-printed with PLA plastic, and its diameter is cm; see Fig. 3. As seen in Fig. 3(a)-(d) and Fig. 3(e)-(h), the robots demonstrate a pronounced condensation at the barrier for both considered center of inertia locations, with characteristic condensation times s for -robots and s for -robots, respectively, see Supplementary Video 1. These values are obtained by averaging over experiments for each type of the robots starting from a randomly arranged swarm. The detailed results are provided in Table 1. In experiments, the measured value of condensation time corresponds to the time frame when all robots localize at the barrier for the first time, with the only allowed exception of one or two robots roaming along closed circular trajectories in the bulk due to the chirality that is inevitably present in their motion.
Once condensed at the boundary, the robots cannot return to the center, which hinders the further emergence of large clusters in the bulk. Moreover, the robots located at the barrier touch it with their noses or lateral surfaces; hence, an axially symmetric micelle cannot form at the barrier. Similar condensation has already been reported Giomi et al. (2013) and even applied to engineer specific swarm behavior Deblais et al. (2018) or collective functionalities Boudet et al. (2021). Its appearance appears absolutely intuitive and is related to the negative curvature of the boundary. As shown in Subsections IV.2 and IV.3, the characteristic times of a single micelle formation are much longer than the condensation times observed in Fig. 3.
@@ IV.1 Boundary condensation of non-chiral robots on a flat surface surrounded by a vertical wall
Chiral trajectories of robots. If each robot moves along closed circular trajectories with diameter much larger than the characteristic robot size , yet much lower than the barrier diameter , the robots can be considered (locally) self-propelling forward, yet they can localize in the bulk and touch the barrier only as a result of collisions with other robots. Moreover, the robot can still return to the bulk after condensing at the barrier if its motion direction along the barrier coincides with its own chirality. Such chiral systems are considered in Refs. Deblais et al. (2018); Barois et al. (2020); Boudet et al. (2021); Siebers et al. (2023).
@@
In the following, we consider two strategies: chiral robots moving along circular trajectories in clockwise or counter-clockwise directions that allow us to demonstrate the emergence of micellization in robotic swarms on flat surfaces, and non-chiral robots in a parabolic potential.
IV.2 Micellization in swarms of chiral robots on a flat surface
@@
We start with considering swarms with an even number of robots in which half of the robots have clockwise (CW) chirality, i.e. move along the circular trajectories centered to the right from the robot’s propulsion vector direction, while the rest half move in the opposite, counter-clockwise (CCW) direction, Fig. 4(a). The robots are made chiral by adding an additional inclination of the bristles in the planes of their arrays. We select the opposite additional inclination angles for the front and bottom bristle arrays, thus adding the rotational component to the robots’ propulsion. In experiments, we set the number of robots rotating clockwise and counter-clockwise equal to zero out the net chirality of the system and avoid chirality-assisted effects such as the formation of unidirectional boundary flows Yang et al. (2020); Barois et al. (2020). As outlined in Sec. IV.1, the robots’ trajectories are tuned in such a way that the radius of the trajectory along which the robot moves is less than the radius of the barrier, but is larger than the size of the robot, which allows us to consider the robot’s motion locally linear. Moreover, the chirality of the robot’s trajectory allows it to turn away from the barrier in certain situations, and thus additionally prevents the condensation mentioned in Sec. IV.1. The considered chiral robots are placed on a flat surface made of plywood and are bounded by a plastic barrier with diameter cm.
@@
In our experiments, we vary the location of the center of inertia considering two cases: -robots [Fig. 4(b)-(e)] and -robots [Fig. 4(f)-(i)]. We capture the collective dynamics of the robotic swarm on video for five minutes and repeat the experiment times for each robot type, considering different randomly chosen initial locations of the robots. In all experiments, we set the high robot activity ().
@@
As seen in Fig. 4(b)-(e), for -robots micelle formation is not observed, although in Fig. 4(c) it is clearly seen that the robots turn away from the boundaries of the barrier without condensing on it (see the red and blue tracks highlighting the trajectories of CW- and CCW-chiral robots, respectively). In contrast, for -robots, a micelle is formed at s, Fig. 4(h), However, is decomposes in several seconds after the formation and is not present in Fig. 4(i), suggesting general instability of micelles in the system with chiral robots. Both cases are illustrated in Supplementary Video 2.
@@
The reason behind such an instability is the following. In a two-robot cluster (representing a quarter of a micelle) composed of a CW-chiral robot placed to the left from a CCW-chiral robot (looking towards the direction of its propulsion vector), two such robots possess propulsion components that attract them together while they propel forward in the same direction. As a result, such a cluster retains its shape during motion. The formation of this type of two-robot cluster is analyzed in more detail in Sec. V.2. However, if we instead put the CW-robot to the right of the CCW-robot, then repulsive propulsion components appear that pull the robots apart. Therefore, such a cluster cannot exist for a long time. The stability of analogous configuration with two CW- or two CCW-chiral robots will, obviously, be intermediate between the two considered cases. The same logic applies to configurations with larger numbers of robots, i.e., half-micelles, complete micelles, and other clusters. In addition, every cluster can be composed with several different combinations of CW- and CCW-chiral robots, for example, the micelle in Fig. 4(h) is composed from CW- and CCW-chiral robots. The number of these combinations increases with the number of robots in the cluster, as is known from combinatorics, and only part of these configurations are stable.
@@
Next, we analyze the statistical properties of the micelles observed in the discussed experiments with swarms of chiral robots on a flat surface. As seen in Table 2, no micelles were observed in the system with -robots. However, in the case of -robots, a micelle is formed in of the experiments, with the mean lifetime s limited by the instability of the micelles formed by chiral robots. The insets of Table 2 highlight CW-chiral robots with red and CCW-chiral robots with blue, showing different ratios of CW- and CCW-chiral robots in the observed micelles. Finally, in accordance with our initial considerations, the characteristic time of micelle formation s significantly exceeds the typical condensation times s (for -robots, see Section IV.1) significantly, justifying the necessity of considering chiral trajectories.
@@
To summarize, condensation at the vertical wall barrier is significantly reduced in swarms of chiral robots moving on a flat surface compared to non-chiral robots. Moreover, micelle formation is observed only in the case of -robots. However, a quantitative analysis of the micellization is complicated by the short micelle lifetime of just a few seconds, making it necessary to consider various strategies toward its increase.
IV.3 Micellization in swarms of non-chiral robots in a parabolic potential

### v3

@@ Abstract
Active matter composed of self-propelled particles features fascinating self-organization phenomena, spanning from motility-induced phase separation to phototaxis to topological excitations depending on the nature and parameters of the system. In the present paper, we consider micelle formation by active particles with a broken symmetry having a circular back and a sharpened nose toward which the particles accelerate. As we demonstrate in experiments with robotic swarms, such particles can either remain in the isotropic phase or form micelles depending on the location of their center of inertia, in accordance with a recent theoretical proposal [T. Kruglov and A. Borisov, Presentations and Videos to 7th Edition of the International Conference on Particle-based Methods (2021), Vol. CT07, p. 2]. Such a behavior is observed for both nonchiral particles moving linearly and placed in a parabolic potential and for chiral particles moving along circular trajectories on a flat surface. By performing experiments with single robots and two-robot collisions, we unveil that the observed emergence of micellization associated with shifting robots’ center of inertia towards their noses is governed by at least two-particle effects, in particular, by a difference in the formation of stable two-robot clusters. Finally, we consider the dependence of micelle lifetime and formation probability as well as two-robot collisions on friction between the lateral surfaces of the robots. Crucially, the predicted micellization does not involve any solvation shells that give rise to the micellization of surfactants but is instead driven by an interplay of activity and particle shape asymmetry.
@@ I Introduction
However, the formation of micelles—round or spherical-shaped assemblies of elongated particles in which they orient one of their nonequivalent edges to the inner region of a cluster and the other one to the outer region—has been demonstrated only in Janus particles covered with surfactants Zhang et al. (2015). This phenomenon appeared elusive in self-propelled particles without charge displacement, until the recent theoretical proposal of active - matter micellization driven purely by particle shape asymmetry Kruglov and Borisov (2021) and experimental studies of rotelles, the rotational analogs of micelles formed by surfactant-like chain of self-rotating active particles Scholz et al. (2021).
@@
In the present paper, we consider ensembles of self-propelled particles that have the teardrop shape shown in Fig. 1(a). Particles having a similar shape occur in biological systems, for example, specific types of blood cells Gütgemann et al. (2014), and can be fabricated on the millimeter to micrometer scale with the help of various techniques, e.g., self-propelled asymmetric Janus particles in Refs. Dai et al. (2016); Shah et al. (2021); Liu et al. (2022). To demonstrate experimentally that such particles can form micellelike clusters depending on the center of inertia location at the particle axis, we construct a swarm of teardrop-shaped vibrating robots (bristle-bots) shown in Figs. 1(b) and 1(c) which are based on the Swarmodroid 1.0 platform Dmitriev et al. (2023). Such robots move with a controlled level of activity and are placed in a shallow parabolic potential of a satellite dish to prevent condensation at the system boundary, typical of self-propelled particles Giomi et al. (2013); Deblais et al. (2018); Boudet et al. (2021), Fig. 1(a), or on a flat surface with their trajectories modified to be slightly chiral (corresponding to robots moving along circular loops). In accordance with the predictions of Ref. Kruglov and Borisov (2021), we report micelle formation both for swarms on a flat surface and in a parabolic potential and address the dependence of their formation probability on the packing density of robots, their center of inertia location, velocity of their motion, and the friction coefficient between lateral surfaces of robots. Finally, we uncover the microscopic mechanisms responsible for such a sharp dependence of the micellization on the location of individual robots’ center of inertia by performing experiments with single robots as well as with two-robot collisions in different conditions.
@@
The paper is organized as follows. In Sec. II, we describe our experimental platform. Then, in Sec. III a key quantitative parameter is introduced which we use to detect micellization. Section IV presents the results of experiments with robotic swarms and includes three subsections. Section IV.1 includes studies of the condensation of robots at the boundary when they propel along straight trajectories on a flat surface. The next Sec. IV.2 features a demonstration of robots’ micellization on a flat surface for robots moving along chiral trajectories and includes the study of the micellization statistical properties. Then, Sec. IV.3 considers micellization of robots in a parabolic potential, including the dependence of the lifetime and formation probability of micelles on the packing density of robots, their motion velocity, and the friction between their lateral surfaces for different locations of the robots’ center of inertia. The experiments considering microscopic origin of the observed micellization transition are discussed in Sec. V. Section V.1 includes studies of individual robot motion in a parabolic potential for different center of inertia locations demonstrating that the behavior at a single-robot level does not demonstrate any pronounced changes. Section V.2 features the studies of two-robot collisions pointing toward a considerable increase in the formation of stable quarter-micelle clusters, hinting towards its possible influence on the emergence of the micellization. Section VI contains the final remarks and discussion of the obtained results.
@@ II Experimental platform for active matter: asymmetric robots
In our experiments, we implement the teardrop-shaped particles as self-propelled bristle-bots converting their vibration to a directed motion, Figs. 1(b)- 1(e). Such robots consist of a printed circuit board [Fig. 1(d)] carrying a vibration motor, a battery, and circuitry for infrared remote control, which allows turning the robots on and off simultaneously and varying their vibration activity (i.e., a self-propulsion velocity) by changing the pulse width modulation (PWM) duty cycle of the motor voltage, and a three-dimensional (3D) printed body with elastic bristles at the bottom [Fig. 1(e)] Dmitriev et al. (2023). The system dynamics is extracted using a tracking pipeline based on our custom recognition software Dmitriev et al. (2023) and the use of ArUco markers.
The length of each robot is , the diameter of the round part is , and the nose angle is corresponding to particles in a complete micelle. We choose such a value of to clearly distinguish micellization from crystallization, as the point symmetry is incompatible with a crystalline order. The height of the robots including the bristles is mm. The robot’s center of inertia lies on its mirror symmetry axis at the distance from its nose cusp. With the help of an additional load, we are able to move the center of inertia from mm (point ), which is the center of inertia of a robot without additional load [Fig. 1(b)], to mm (point ) for a robot with a load [Fig. 1(c)]. We attach a load consisting of an M616 screw and three M6 nuts made of steel, which are fastened to a hole near the robot’s nose, to move the center of inertia to point . In the following, we denote such robots as -robots. To increase the mass of a robot by the same amount while maintaining the center of inertia at point , we instead add a load consisting of a steel DIN125A M18 washer, which is glued to the robot cap at point Sup . We denote these robots as -robots. The mass of -robots and -robots is the same and equal to g. We study swarms composed of , , , and robots.
Figures 1(f) and 1(g) show the velocities of - and -robots, respectively, as dependencies of the PWM duty cycle, which represents the vibration activity. Due to the complex dynamics involved in the conversion of vibration to directional motion by bristles Giomi et al. (2013); Scholz et al. (2016); Porvatov et al. (2021); Dmitriev et al. (2023), there is a distribution of velocities between different robots at each value of the vibration activity, which we show as a color-encoded histogram. The standard deviation of this distribution reaches one third of the mean value for -robots at the vibration activity of . Therefore, we do not refer directly to the robot velocity throughout the paper, but instead use vibration activity which is exactly reproducible. It is seen that, despite their different weight distributions, the velocities of -robots and -robots coincide for vibration activities up to that are studied in the following.
@@ III Micellization parameter
where and are the in-plane coordinates of the noses of particles with indices and , is the number of particles in a complete micelle (in our case, ), is the characteristic distance between the noses of particles at which they are considered forming a micelle (here, we choose mm), and is the total number of particles in the system. Figure 2(a) shows the values of such micellization parameter extracted experimentally for several model systems: an isolated complete micelle with , several types of isolated incomplete micelles with , an inverse micelle with , a structure of densely packed robots with crystalline order, a single robot, and two types of two-robot clusters (“tandem” and “heart”) that, as will be shown in Sec. V, are important for understanding the processes facilitating micelle formation. The terms in the sum by index tend to unity for a complete micelle while rendering values orders of magnitude smaller for other configurations. However, due to the exponential dependence on the distance between the robots’ noses, the micellization parameter yields values somewhat lower than unity for experimentally observed micelles, with the majority of observed micelles corresponding to . For incomplete micelles and nonmicellar structures, the micellization parameter yields exponentially low values. The normalization prefactor ensures the invariance of the parameter with respect to the number of robots. For fully chaotic systems, the micellization parameter yields the value independently of the number of robots in the system. The same value corresponds to the system consisting of a single robot and is the lowest possible value of the micellization parameter.
We also study the dependence of the micellization parameter on the distance between the particles in a single micelle by introducing random shifts of the particles’ noses from the center of a micelle, Fig. 2(b). The components of the shift vectors are normally distributed with the zero mean and the dispersion that varies from to mm (approximately of the robot’s diameter). The micellization parameter decreases exponentially depending on this amplitude, producing the same value for a complete micelle with mm gaps between the robots and an ideal incomplete micelle without one robot.
IV Experimental studies of micellization in robotic swarms
@@ IV.1 Boundary condensation of non-chiral robots on a flat surface surrounded by a vertical wall
We start with considering a swarm of teardrop-shaped robots moving on a flat surface and confined by a circular barrier (Table 1). The barrier is 3D printed with PLA plastic, and its diameter is cm; see Fig. 3. As seen in Figs. 3(a)-3(d) and Fig. 3(e)-3(h), the robots demonstrate a pronounced condensation at the barrier for both considered center of inertia locations, with characteristic condensation times s for -robots and s for -robots, respectively, see supplementary video 1. These values are obtained by averaging over experiments for each type of the robots starting from a randomly arranged swarm. The detailed results are provided in Table 1. In experiments, the measured value of condensation time corresponds to the time frame when all robots localize at the barrier for the first time, with the only allowed exception of one or two robots roaming along closed circular trajectories in the bulk due to the chirality that is inevitably present in their motion.
Once condensed at the boundary, the robots cannot return to the center, which hinders the further emergence of large clusters in the bulk. Moreover, the robots located at the barrier touch it with their noses or lateral surfaces; hence, an axially symmetric micelle cannot form at the barrier. Similar condensation has already been reported Giomi et al. (2013) and even applied to engineer specific swarm behavior Deblais et al. (2018) or collective functionalities Boudet et al. (2021). Its appearance appears absolutely intuitive and is related to the negative curvature of the boundary. As shown in Secs. IV.2 and IV.3, the characteristic times of a single micelle formation are much longer than the condensation times observed in Fig. 3.
@@
Chiral trajectories of robots. If each robot moves along closed circular trajectories with diameter much larger than the characteristic robot size , yet much lower than the barrier diameter , then the robots can be considered (locally) self-propelling forward, yet they can localize in the bulk and touch the barrier only as a result of collisions with other robots. Moreover, the robot can still return to the bulk after condensing at the barrier if its motion direction along the barrier coincides with its own chirality. Such chiral systems are considered in Refs. Deblais et al. (2018); Barois et al. (2020); Boudet et al. (2021); Siebers et al. (2023).
@@
In the following, we consider two strategies: chiral robots moving along circular trajectories in clockwise or counterclockwise directions that allow us to demonstrate the emergence of micellization in robotic swarms on flat surfaces, and non-chiral robots in a parabolic potential.
IV.2 Micelle-like clusters in swarms of chiral robots on a flat surface
@@
We start with considering swarms with an even number of robots in which half of the robots have clockwise (CW) chirality, i.e., move along the circular trajectories centered to the right from the robot’s propulsion vector direction, while the rest half robots move in the opposite, counterclockwise (CCW) direction, Fig. 4(a). The robots are made chiral by adding an additional inclination of the bristles in the planes of their arrays. We select the opposite additional inclination angles for the front and bottom bristle arrays, thus adding the rotational component to the robots’ propulsion. In experiments, we set the number of robots rotating clockwise and counterclockwise equal to zero out the net chirality of the system and avoid chirality-assisted effects such as the formation of unidirectional boundary flows Yang et al. (2020); Barois et al. (2020). As outlined in Sec. IV.1, the robots’ trajectories are tuned in such a way that the radius of the trajectory along which the robot moves is less than the radius of the barrier, but is larger than the size of the robot, which allows us to consider the robot’s motion locally linear. Moreover, the chirality of the robot’s trajectory allows it to turn away from the barrier in certain situations, and thus additionally prevents the condensation mentioned in Sec. IV.1. The considered chiral robots are placed on a flat surface made of plywood and are bounded by a plastic barrier with diameter cm.
@@
In our experiments, we vary the location of the center of inertia considering two cases: -robots [Figs. 4(b)-reffig:Chiral(e)] and -robots [Figs. 4(f)-4(i)]. We capture the collective dynamics of the robotic swarm on video for five minutes and repeat the experiment times for each robot type, considering different randomly chosen initial locations of the robots. In all experiments, we set the high robot activity ().
@@
As seen in Figs. 4(b)-4(e), for -robots micelle formation is not observed, although in Fig. 4(c) it is clearly seen that the robots turn away from the boundaries of the barrier without condensing on it (see the red and blue tracks highlighting the trajectories of CW- and CCW-chiral robots, respectively). In contrast, for -robots, a micelle is formed at s, Fig. 4(h). However, it decomposes in several seconds after the formation and is not present in Fig. 4(i), suggesting general instability of micelles in the system with chiral robots. Both cases are illustrated in supplementary video 2.
@@
The reason behind such an instability is the following. In a two-robot cluster (representing a quarter of a micelle) composed of a CW-chiral robot placed to the left from a CCW-chiral robot (looking towards the direction of its propulsion vector), two such robots possess propulsion components that attract them together while they propel forward in the same direction. As a result, such a cluster retains its shape during motion. The formation of this type of two-robot cluster is analyzed in more detail in Sec. V.2. However, if we instead put the CW-robot to the right of the CCW-robot, then repulsive propulsion components appear that pull the robots apart. Therefore, such a cluster cannot exist for a long time. The stability of analogous configuration with two CW- or two CCW-chiral robots will, obviously, be intermediate between the two considered cases. The same logic applies to configurations with larger numbers of robots, i.e., half-micelles, complete micelles, and other clusters. In addition, every cluster can be composed with several different combinations of CW- and CCW-chiral robots, for example, the micelle in Fig. 4(h) is composed from three CW- and five CCW-chiral robots. The number of these combinations increases with the number of robots in the cluster, as is known from combinatorics, and only part of these configurations are stable.
@@
Next, we analyze the statistical properties of the micelles observed in the discussed experiments with swarms of chiral robots on a flat surface. As seen in Table 2, no micelles were observed in the system with -robots. However, in the case of -robots, a micelle is formed in of the experiments, with the mean lifetime s limited by the instability of the micelles formed by chiral robots. The insets of Table 2 highlight CW-chiral robots with red and CCW-chiral robots with blue, showing different ratios of CW- and CCW-chiral robots in the observed micelles. Finally, in accordance with our initial considerations, the characteristic time of micelle formation s significantly exceeds the typical condensation time s (for -robots, see Sec. IV.1) significantly, justifying the necessity of considering chiral trajectories.
@@
To summarize, condensation at the vertical wall barrier is significantly reduced in swarms of chiral robots moving on a flat surface compared to nonchiral robots. Moreover, micelle formation is observed only in the case of -robots. However, a quantitative analysis of the micellization is complicated by the short micelle lifetime of just a few seconds, making it necessary to consider various strategies toward its increase.
IV.3 Micellelike clusters in swarms of nonchiral robots in a parabolic potential
We proceed with studying a system of nonchiral -robots moving with low activity ()[Fig. 5(a)] and placed in a parabolic satellite dish having the dimensions of [Fig. 1(a)]. In contrast to a hard wall, such a dish creates a soft localizing potential and, as we will demonstrate further, prevents the condensation of robots at the boundary. The micellization parameter fluctuates during the system evolution but does not demonstrate any pronounced growth characteristic of micellization, Fig. 5(a). In this case, robots are predominantly directed away from the vertex of the parabolic potential Dauchot and Démery (2019). Covering the lateral surfaces of robots with abrasive tape in order to increase the friction between them does not result in any systematic changes of the micellization parameter, see the blue solid line in Fig. 5(a).
Next, we consider the dependencies of the micellization parameter on the activity of the robots, their center of inertia location, and the packing density that is varied by changing the number of robots in the parabolic potential. For systems of -robots with low () and high () activity shown in Figs. 6(a) and 6(b) the micellization parameter fluctuates near the minimal level () independently of the density of the system, and micelle formation is not observed; see Table 3. -robots do not demonstrate any micelles as well in the experiments with sparse swarms of robots; see Figs. 6(c) and 6(d). However, stable half-micelles appear in these systems, characterized by an increase in the micellization parameter (). The formation of transient micelles with a short lifetime of a few seconds emerges at the higher density , and is well expressed for , indicated by multiple peaks of the micellization parameter in Figs. 6(c) and 6(d) both for low () and high () activities. Increasing robot activity results in shorter average micelle lifetimes, namely s for and s for for high robot activity, compared to s for and s for for low activity.
Next, we modify the system by covering the lateral surfaces of all robots with an abrasive (P800 sandpaper) Sup to consider how the micellization changes in the case of high friction between the particles, which may prove important for microscale implementations. As seen in Figs. 6(e) and 6(f), the behavior of -robots is similar to the case of particles with low lateral friction; Figs. 6(a) and 6(b). The micellization parameter slightly fluctuates, indicating the absence of micellization. The same is confirmed by the data in Table 3.